OpenStudy (anonymous):

Rewrite the rational exponent as a radical. 5 to the 3 over 4 power, to the 2 over 3 power

OpenStudy (ian_sage015):

I'm drunk

OpenStudy (anonymous):

ok

OpenStudy (ian_sage015):

bye

OpenStudy (ian_sage015):

goode luck

OpenStudy (anonymous):

I need help on a question can u help me

OpenStudy (ian_sage015):

i can try

OpenStudy (anonymous):

okay thanks

OpenStudy (ian_sage015):

what do you think it is

OpenStudy (anonymous):

I am not sure that is why I asked for help :)

OpenStudy (ian_sage015):

haha tbh I'm confused too

OpenStudy (ian_sage015):

do you know anything about radical

OpenStudy (anonymous):

yea but this one is hard for me

OpenStudy (ian_sage015):

hmm hold on ill try to figure it out

OpenStudy (anonymous):

ok

OpenStudy (ian_sage015):

Rational Exponents. Fractional Exponents. The use of rational numbers as exponents. A rational exponent represents both an integer exponent and an nth root. The root is found in the denominator (like a tree, the root is at the bottom), and the integer exponent is found in the numerator.

OpenStudy (ian_sage015):

so whats the root

OpenStudy (anonymous):

hold on let me see

OpenStudy (faiqraees):

\[\large\rm a^{x^y} = a^{x*y} \]

OpenStudy (ian_sage015):

yes^

OpenStudy (ian_sage015):

nvm im drunk can't focus. bye

OpenStudy (anonymous):

okay bye

OpenStudy (instagrammodel):

(53/4)2/3 = 51/2 = √5